OpenStudy (bananas1234):

Arrange the list of numbers in increasing order. 3/5, 5/9, 4/7

OpenStudy (bananas1234):

3/5 = 0.60000 5/9 = 0.55555 4/7 = 0.57142 So 5/9 < 4/7 < 3/5 is that right?

OpenStudy (bananas1234):

so it would be 5/9 ,4/7, 3/5?

OpenStudy (bananas1234):

@Nnesha

OpenStudy (rational):

Excellent! http://www.wolframalpha.com/input/?i=5%2F9+%3C+4%2F7+%3C+3%2F5

OpenStudy (bananas1234):

Thanks!